# Dissolution of non-Hermitian skin effect in one-dimensional lattices with linearly varying nonreciprocal hopping

### v1

@@ V Summary
In this paper, we study the 1D non-Hermitian lattices with linearly increasing nonreciprocal hopping between the nearest-neighboring sites. We find that due to the spatially varying nonreciprocity, the eigenenergies and eigenstates behave quite differently from those of non-Hermitian systems with constant nonreciprocity. When the nonreciprocity is weak, some of the eigenenergies under OBC remain real and form an equally spaced ladder. The corresponding eigenstates are localized at one end of the 1D lattice due to the NHSE and exhibit a Gaussian distribution. The rest of the eigenenergies are imaginary with the eigenstates shifted from the boundary into the bulk and forming tightly bound states. As the strength of nonreciprocity increases, the energy spectrum undergoes a real-imaginary transition, accompanied by the shifting of eigenstates from the boundary to the bulk. Thus, the NHSE is dissolved gradually. When the nonreciprocity becomes strong enough, there will be no NHSE in the system, and all the states are Gaussian bound states in the bulk. These results reveal the exotic properties of non-Hermitian lattices with spatially varying nonreciprocity and open a door for future studies on such systems.
@@ Acknowledgements.
This work is supported by NSFC (Grant No. 12204326), Beijing Natural Science Foundation (Grant No. 1232030), and R&D Program of Beijing Municipal Education Commission (Grant No. KM202210028017). R.-L is supported by the NSFC under Grant No. 11874234 and the National Key Research and Development Program of China (2018YFA0306504).

### v2

@@ V Summary
In this paper, we study the 1D non-Hermitian lattices with linearly increasing nonreciprocal hopping between the nearest-neighboring sites. We find that due to the spatially varying nonreciprocity, the eigenenergies and eigenstates behave quite differently from those of non-Hermitian systems with constant nonreciprocity. When the nonreciprocity is weak, some of the eigenenergies under OBC remain real and form an equally spaced ladder. The corresponding eigenstates are localized at one end of the 1D lattice due to the NHSE and exhibit a Gaussian distribution. The rest of the eigenenergies are imaginary with the eigenstates shifted from the boundary into the bulk and forming tightly bound states. As the strength of nonreciprocity increases, the energy spectrum undergoes a real-imaginary transition, accompanied by the shifting of eigenstates from the boundary to the bulk. Thus, the NHSE is dissolved gradually. When the nonreciprocity becomes strong enough, there will be no NHSE in the system, and all the states are Gaussian bound states in the bulk. It is quite interesting to see that the increment in nonreciprocal hopping does not necessarily lead to the enhancement of NHSE. On the contrary, the model we studied here shows that the linear variation in the nonreciprocal hopping terms can dissolve the NHSE. Our work reveals the exotic properties of non-Hermitian lattices with spatially varying nonreciprocity and opens a door for future studies on such systems. As to the experimental realization, it has been reported that nonreciprocal hopping can be realized by using photonic coupled resonant optical waveguides, where judicious optical gain and loss elements in the coupling link rings can be designed [77]. The variation in the hopping terms can be realized by tuning the distance between the neighboring waveguides.
@@ Acknowledgements.
This work is supported by the NSFC (Grant No. 12204326), Beijing Natural Science Foundation (Grant No. 1232030), and R&D Program of the Beijing Municipal Education Commission (Grant No. KM202210028017). R.-L. is supported by the NSFC under Grant No. 11874234, the National Key Research and Development Program of China (Grant No. 2018YFA0306504), and the Ministry of Science and Technology of China (Grand No. 2023YFC2205800).